OpenStudy (anonymous):

Determine if the function f(x) = 1 + x − x^3 satisfies the hypotheses of the Mean Value Theorem (MVT) on the interval [−2, 1]. If it does, find all possible values of c

ganeshie8 (ganeshie8):

a polynomial is continuous and differentiable everywhere, given function is a polynomial so it satisfies the hypothesis of MVT

ganeshie8 (ganeshie8):

next, find f'(x)

OpenStudy (dumbcow):

\[f'(c) = \frac{f(1) -f(-2)}{1-(-2)}\] \[f'(c) = -2\] \[1-3c^{2} = -2\] \[c = \pm 1\]